A teacher asked pupils to divide 28 by 3
Gail wrote : 28 ÷ 3 = 9.33
Ahmed wrote : 28 ÷ 3 = 9 1/3
The teacher marked both correct, but said Ahmed’s answer was better than Gail’s answer. Explain why Ahmed’s answer is not the same as Gill’s answer
1
Expert's answer
2020-10-12T17:46:05-0400

Ahmed is right. His answer is entirely correct


"\dfrac{28}{3}=\dfrac{27+1}{3}=\dfrac{27}{3}+\dfrac{1}{3}=9+\dfrac{1}{3}=9\dfrac{1}{3}"

Gail is not right. Why does he round up the result to 2 decimal places?


"\dfrac{28}{3}=9.333333333...=9.(3)\approx9.33"

Does the teacher say to round up the result of dividing? It is unknown.

I prefer the Ahmed’s answer.